STUDY OF CONFORMATIONAL LANDSCAPE AND TUNNELING MOTIONS IN 4-FLUOROBENZYL ALCOHOL^{a}

\underline{ALEXEI E. NIKOLAEV}, DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

High resolution UV study of 4-fluorobenzyl alcohol (FBA) was reported previously in 2001 ^{a}and some preliminary analysis was performed. We performed additional extensive conformer search calculations and nuclear spin statistics analysis to explain peculiar splittings and intensity variations of FBA rotationally resolved electronic spectra. Interpretation of the structural based on these results will be discussed.

aWork supported by NSF.

aA.E.Nikolaev and D.W.Pratt 56th Int.Symp. Molec. Spectrosc. Paper RF03, June 2001